# Conceptualization of this study, Methodology, Writing

### v1

@@ 3 Proposed Framework
The action segmentation task involves predicting the action category for each frame in a video. Formally, given a sequence of $T$-frame features $X=[\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{T}]\in\mathbb{R}^{T\times d},\quad\text{where }\mathbf{x}_{t}\in\mathbb{R}^{d}$ represents the $d$-dimensional feature of the $t$-th frame. The objective is to predict a sequence of actions $\hat{Y}=[\hat{y}_{1},\hat{y}_{2},\ldots,\hat{y}_{T}],\quad\text{where }\hat{y}_{t}\in C.$ The set $C=\{0,1,2,\ldots,C,-100\}$ includes all action categories in the dataset, with $-100$ representing the background class. The frame-wise ground truth labels are $Y=[y_{1},y_{2},\ldots,y_{T}],$ accompanied by segment information $S=\{(y_{s1},y_{e1},a_{c1}),\ldots,(y_{sn},y_{en},a_{cn})\},$ where $y_{s}$ and $y_{e}$ denote the start and end of a segment, $a_{c}$ represents the action class label, and $n$ is the number of segments, including background segments and notations are listed in Table 1. The goal is to accurately predict action transitions and segment boundaries. Since action transitions are continuous and boundaries are often subjective, considering this 5% of the total segment length is labeled as a transition buffer zone at the start and end of each segment.
Our proposed framework is depicted in figure 3 that enables three distinct parts. First, the frame wise feature extraction, second, Encoder-Decoder architecture with hierarchical sliding window attention, and last, boundary aware loss function that guides to detect the transition with start-center-end points.
Figure 3: Proposed Multi-Stage Boundary-Aware Transformer Network (MSBATN) architecture
@@ 3.1 Frame-Wise Feature Extraction
Dino with Register Oquab et al. (2023) (Dinov2-Reg) captures complex spatial features in a self-supervised manner, Figure 4 depicts the performance in comparison with ViT and Resnet.
@@
Dinov2-Reg model extracts features using a self-supervised vision transformer model, which processes image-like inputs (frames in this case) to output dense, high-dimensional feature representations. The input feature dimensions (256, 1536), where 256 denotes the 16x16 patch dimension, and 1536 is the feature dimension. A 2D convolutional layer with a kernel size of (256, 1536) is applied to the input frames, which projects the high-dimensional DINO features (256x1536) into a lower-dimensional representation (1x2048). This is equivalent to performing dimensionality reduction via a convolutional operation, which retains important information while reducing the computational complexity for further processing in a Multi-Stage Boundary-Aware Transformer Network (MSBATN). For more details, please refer to the attached supplementary material file.
@@ 3.2 Multi-Stage Boundary-Aware Transformer Network (MSBATN)
We present MSBATN transformer-based encoder-decoder architecture without positional encoding to improve the performance of the segmentation task. Each encoder-decoder block comprises a series of TCN and hierarchical sliding window attention modules.
@@ 3.2.1 Encoder-Decoder
The frame-wise features $(T,D)$ are passed through an encoder block. The encoder block comprises a series of dilated Acausal Temporal Convolution feed-forward layer blocks to capture context from past and future, followed by the hierarchical sliding window attention module and, lastly, project the encoded features. Instead of using traditional fully connected layers, dilated temporal convolutions are employed to expand the receptive field without increasing the number of parameters or sacrificing resolution. It introduces local inductive biases to effectively capture temporal dependencies and patterns over extended sequences. The model adopts a hierarchical pattern in which the attention layers simultaneously focus on local-global information, mirroring techniques utilized in longformerBeltagy et al. (2020). To capture the global context, we calculated the hierarchical attention at different scales, and then we aggregated to obtain the global context. A detailed discussion is added in the next subsection.
@@
The decoder considered the temporal relationships between action-boundary segments to improve the initial predictions. Similar to the encoder block, the decoder begins with a temporal convolutions block, attention module and output. However, instead of acausal convolution, it has causal convolution, so it does not have any information about the future. Each attention block uses a weighted residual connection to combine the outputs from the feed-forward layers, with the weight $\alpha$ decreasing exponentially across decoders to prevent error accumulation. Last, the decoder generates action boundary labels. The network is comprised of one encoder and three decoders, with each encoder and decoder containing ten TCN-attention blocks. We set the feature dimension in the encoder and decoder layer to 256. Additionally, a temporal dropout layer is applied to the input feature of the encoder, which randomly drops the entire feature channel with a dropout rate of 0.3 to mitigate the over-fitting issue.
3.2.2 Hierarchical dual Sliding Window Attention
Instead of computing attention scores across all token pairs, as in self-attention, we employed a sparse-attention mechanism. This approach uses a hierarchical sliding window that restricts attention to a fixed-size window around each token at a specific layer. The window slides and aggregates scores, meaning each token attends only to its neighboring tokens within the specified window. This reduces both space and computational complexity to $O(n)$. To capture temporal dependencies at varying distances, a dilated attention window is applied with increasing dilation rates $(0,1,2,3,\dots)$ in each layer. This approach allows for capturing both local details from individual frames and long-term temporal patterns across the sequence. The action head and boundary head predict the action and start-end labels, respectively Nwoye et al. (2022). For a sequence length $T$ and we define the base temporal window size $64$, smallest window size $16$, largest window size $512$, average number of action segments $S$. To expand the receptive field, the effective window size $W_{d}$ for each layer $d$ is calculated using dilation as $W_{d}=W\cdot(1+d)$ where $W$ is the base window size, $d$ is the dilation factor ($d=0,1,2,\dots$). For Dual Window Sliding Attention, the effective receptive field starts at $16$ and grows up to $512$, that means $W_{d}\in[16,512]$. Local attention is computed within each window and then aggregated across windows. The temporal hierarchical attention operates at multiple scales. The scale $T_{s}$ is reduced iteratively $T_{s}=\frac{T}{2^{s}}$ for $s=0,1,2,\dots$, until $T_{s}\geq 512$. The average number of scales is approximately $\log_{2}\left(\frac{T}{S}\right)$, based on the sequence length $T$ and the average number of action segments $S$. At each scale $T_{s}$, global attention weights are computed using a sparse mechanism with dilation. For a token $x_{i}$ attending to token $x_{j}$ at scale $T_{s}$, the attention weight $\alpha_{ij}^{s}$ is
$\alpha_{ij}^{s}=\frac{\exp\left(e_{ij}^{s}\right)}{\sum_{k\in\mathcal{N}_{i}^{s}}\exp\left(e_{ik}^{s}\right)}$ (1)
where $e_{ij}^{s}$ is the attention score at scale $T_{s}$, $\mathcal{N}_{i}^{s}$ is the neighborhood of token $x_{i}$ at scale $T_{s}$, defined by the dilated window size $W_{d}$. The final attention is calculated by aggregating scores across all scales and normalizing them with a softmax function:
$\alpha_{ij}=\frac{\exp\left(\sum_{s}e_{ij}^{s}\right)}{\sum_{k\in\mathcal{N}_{i}}\exp\left(\sum_{s}e_{ik}^{s}\right)}$ (2)
where $\mathcal{N}_{i}=\bigcup_{s\in S}\mathcal{N}_{i}^{s}$ is the union of neighborhoods across all scales, $\alpha_{ij}$ is the final attention weight for token $x_{i}$ attending to $x_{j}$. At each scale $T_{s}$, the attention score $e_{ij}^{s}$ between tokens $x_{i}$ and $x_{j}$ within the neighborhood $\mathcal{N}_{i}^{s}$ is:
$e_{ij}^{s}=\text{sim}(x_{i},x_{j}^{s})$ (3)
where $x_{j}^{s}$ is the representation of token $x_{j}$ at scale $s$, $\text{sim}(\cdot,\cdot)$ is a similarity function (e.g., dot product). By restricting computations to sparse attention neighborhoods $\mathcal{N}_{i}$, the computational overhead is significantly reduced while retaining efficiency. The number of scales is determined byNumber of $\text{Number of scales}=\log_{2}\left(\frac{T}{S}\right).$. The scale reduction follows $T_{s}=\frac{T}{2^{s}},\quad s=0,1,2,\dots,\log_{2}\left(\frac{T}{S}\right),$ reducing the sequence length $T$ to approximately $\frac{T}{S}$. For example, if $T=1000$ and $S=4$, the segment sizes would be $(1000,500,250,125,\dots)$.
@@ 3.3 Boundary-aware Segmentation Loss
In temporal action segmentation, accurately capturing action transitions and segment-level dynamics across varying time scales is crucial. Surgical videos are typically long, and the duration of the segment fluctuates significantly, as shown in Table 2. To address this, we propose a Boundary-Aware Segmentation Loss that integrates multiple loss functions, each targeting specific aspects of the segmentation task. To handle class imbalances and ensure precise classification at each timestep, we employ Focal Loss. This loss function down-weights well-classified examples, allowing the model to focus on harder, misclassified examples. This approach is particularly beneficial in datasets where certain action classes are underrepresented. To enhance the overlap between predicted and ground truth segments, we utilize Dice Loss derived from the Dice Coefficient:
@@
This formulation emphasizes the overlap between predictions and ground truth, making it sensitive to both false positives and false negatives. However, focal and dice loss primarily focuses on frame-wise classification accuracy without explicitly considering the temporal relationships between frames. As can be seen in figure 5, the uppermost figure shows the similarity between neighbouring frames, which are ambiguous; we smoothen the distribution by measuring centre-boundary Gaussian weighted cosine similarity depicted in the middle of the figure. To capture temporal dependencies and ensure smooth transitions within action segments, we introduce a loss term based on Gaussian-weighted cosine similarity. This loss penalizes deviations in feature representations between consecutive frames, weighted by a Gaussian function centred on action segments:
$\mathcal{L}_{\text{sim}}=\sum_{t=1}^{T-1}G(t,t+1)\left(1-\cos(\mathbf{f}_{t},\mathbf{f}_{t+1})\right)$ (6)
where $G(t,t+1)$ is the Gaussian weight emphasizing temporal proximity, and $\cos(\mathbf{f}_{t},\mathbf{f}_{t+1})$ denotes the cosine similarity between feature vectors at frames $t$ and $t+1$. To further refine boundary precision, we incorporate a Gaussian Similarity Truncated Mean Squared Error (MSE) loss. This loss focuses on boundary regions, applying a Gaussian weight to the MSE, which emphasizes errors near segment boundaries:
@@
Where $G_{\text{boundary}}(t)$ is a Gaussian weight centred on the boundary, $\hat{b}_{t}$ and $b_{t}$ are the predicted and ground truth boundary positions at frame $t$, and $\tau$ is a truncation threshold to limit the influence of large errors. By integrating these loss components, the total temporal loss function is defined as:
@@
Where $\alpha$, $\beta$, $\gamma$, and $\delta$ are weighting factors that balance the contributions of each loss term. This comprehensive loss function enables the model to effectively capture class imbalances, segment overlaps, temporal dependencies, and boundary precision, leading to improved performance in temporal action segmentation tasks.
@@ 4 Experiments and Analysis
Table 3: Comparison of Models Across Datasets (SAR-RARP, Jigsaw, Cholec80)

### v2

@@ 3 Proposed Framework
The action segmentation task involves predicting the action category for each frame in a video. Formally, given a sequence of $T$-frame features $X=[\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{T}]\in\mathbb{R}^{T\times d},\quad\text{where }\mathbf{x}_{T}\in\mathbb{R}^{d}$ represents the $d$-dimensional feature of the $t$-th frame. The objective is to predict a sequence of actions $\hat{Y}=[\hat{y}_{1},\hat{y}_{2},\ldots,\hat{y}_{T}],\quad\text{where }\hat{y}_{T}\in C.$ The set $C=\{0,1,2,\ldots,c\}$ includes all action categories in the dataset. The frame-wise ground truth labels are $Y=[y_{1},y_{2},\ldots,y_{T}],$ accompanied by segment information $S=\{(y_{s_{1}},y_{e_{1}},a_{c_{1}}),\ldots,(y_{s_{n}},y_{e_{n}},a_{c_{n}})\},$ where $y_{s}$ and $y_{e}$ denote the start and end of a segment, $a_{c}$ represents the action class label, and $n$ is the number of segments, including background segments. The goal is to accurately predict action transitions and segment boundaries. Since action transitions are continuous and boundaries are often subjective, considering this 5% of the total segment length is labeled as a transition buffer zone at the start and end of each segment.
Our proposed framework, the Multi-Stage Boundary-Aware Transformer Network (MSBATN), is illustrated in Figure 3. MSBATN is designed to improve action segmentation in untrimmed surgical videos by accurately identifying action boundaries, even in lengthy and complex sequences. It achieves this through three main components: robust frame-wise feature extraction, an encoder-decoder architecture based on TCN-Transformer-TCN that incorporates Hierarchical Sliding Window Attention (HSWA) and specialised acausal/causal convolutions, and a novel boundary-aware loss function.
Figure 3: Proposed Multi-Stage Boundary-Aware Transformer Network (MSBATN): Input video features $(X)$ are first processed by TCN encoder consisting Acausal Conv1D (see right bottom) for temporal feature extraction to produce $X_{tcn}$. These features are then fed into the Transformer Encoder, consisting of Hierarchical Sliding-Window Attention (HSWA) (see right top), to generate $X_{encoded}$. The $X_{encoded}$ features are expanded to the original temporal resolution to separate Action and Boundary segmentation heads. Finally, a TCN Decoder refines these initial predictions to output the final action segments and boundaries.
@@ 3.1 Frame-Wise Feature Extraction
The pipeline begins with an input surgical video from which individual frames are extracted. For MSBATN, we specifically use Dinov2-Reg, a feature extractor that improves upon standard ViTs. In standard ViTs, images are divided into patches represented by tokens, but "outlier tokens" can introduce errors due to unusually high values. Dinov2-Reg addresses this by adding register tokens as dedicated placeholders, enhancing feature robustness and reducing reliance on outlier tokens. As shown in Figure 4, Dinov2-Reg Oquab et al. (2023) captures complex spatial features more effectively in a self-supervised manner compared to standard ViT and ResNet methods.
@@
The Dinov2-Reg extractor converts each frame into a high-dimensional patch-based feature representation (e.g., resulting in a ($H^{\prime}\times W^{\prime}\times C_{dino}$) tensor. We use (256, 1536), where 256 denotes the $16\times 16$ patch dimension, and 1536 is the feature dimension. To create a fixed-size feature vector per frame suitable for our temporal model, we project these features. Specifically, we apply a 2D convolutional layer (or an equivalent linear projection on flattened patch features) to transform the DINOv2 output for each frame into a (d)-dimensional vector (d=2048). This results in an initial feature sequence X with dimensions (T, D), where T is the sequence length. This step standardises the frame features and prepares them for the subsequent MSBATN architecture.
@@ 3.2 Multi-Stage Boundary-Aware Transformer Network (MSBATN)
The core of our framework is the MSBATN, a Transformer-inspired encoder-decoder architecture designed for effective temporal feature learning and segmentation (see Figure 3). It includes an initial TCN-based feature processor, a Transformer-based encoder for deeper contextual understanding, and TCN-based decoders for refinement. Notably, MSBATN does not need explicit positional encodings, as its design naturally captures sequence order and relationships through temporal convolutions and attention mechanisms.
@@ 3.2.1 Encoder-Decoder
The encoder path begins by processing the frame-wise features $X=(B,D_{in},T_{orig})$, where B is batch size, $D_{in}$ is the input feature dimension, and $T$ is the original sequence length. The input features first pass through an initial TCN stage. It projects the input dimension (2048) to the model’s working dimension (256). It employs dilated 1D Acausal TCN where the kernel k is centred on the current frame, covering T-K T, T+K. In Causal TCN, the kernel K shifts from right to left, that is, T-K to T thus never peeks into the future, and therefore safe for real-time or streaming inference (Figure 3). When the sequence is too long we applied stride greater than 1 to reduce the sequence length $T$ to $T_{reduced}$ resulting in $X_{tcn}=(B,256,T_{reduced})$. This makes subsequent attention computations more efficient. The dilated convolutions allow the network to build a hierarchical receptive field, capturing local temporal patterns and dependencies over extended sequences effectively. These TCN blocks can access both past and local future context within their receptive field during training, enabling robust local feature extraction.
@@
$X_{tcn}$ is then processed by a stack of encoder blocks. Each block aims to capture more complex and longer-range temporal dependencies. Each encoder block consists of Dual Sliding Window Attention (DSWA), Hierarchical Temporal Attention (HTA) and Feed-Forward Network (MLP) with Layer Normalisation and Residual Connections. Lastly the initial predictions from the encoder are refined by the decoder. Instead of acausal convolution in the encoder TCN, the decoder uses causal convolution to avoid information about the future. The network is comprised of one encoder and three decoders, with each encoder and decoder containing ten TCN blocks. We set the feature dimension in the encoder and decoder layers to 256. Additionally, a temporal dropout layer is applied to the input feature of the encoder, which randomly drops the entire feature channel with a dropout rate of 0.3 to mitigate the over-fitting issue.
3.2.2 Hierarchical Sliding Window Attention
We present a Hierarchical Sliding Window Attention (HSWA) module that effectively captures temporal dependencies by combining dynamic dual-sliding window attention (DSWA) with hierarchical temporal attention for global context. In DSWA, instead of dense self-attention, we use a sparse attention approach inspired by Longformer Beltagy et al. (2020). For each query token (frame representation), attention is computed over two distinct local windows. Across the stack of encoder layers, one set of windows progressively expands (e.g., from a one-sided window of 16 up to 256, for an effective receptive field of 32 to 512), while the other set of windows symmetrically shrinks. This dynamic dual-view allows the model to capture dependencies at varying scales throughout the network depth. Within each sliding window, attention is dilated (e.g., with rates 0, 1, 2, …) in deeper layers. The dilation rates start from 0, meaning some windows, especially in earlier layers or specific attention heads, operate without dilation, focusing squarely on immediate local context. This allows the window to cover a larger receptive field without increasing the number of attended tokens. This dynamically sized windows, ensures that smaller, more focused windows are always present to capture fine-grained local details, even as other windows expand or become more dilated to see broader context. To complement the local/sparse sliding windows and capture broader context, we incorporate a HTA at multiple temporal scales. The sequence length $T_{reduced}$ is conceptually downsampled e.g., $T_{s}=T_{\text{reduced}}/2^{s}$, for scales $(s=0,1,2,\ldots)$. At each scale $(T_{s})$, sparse attention is computed. The number of scales can be determined by $\log_{2}(T_{\text{reduced}}/S_{\text{avg}})$, where $(S_{\text{avg}})$ might relate to an average segment length or a minimum sensible sequence length for global attention (e.g., until $(T_{s}\geq 512)$ as mentioned, though this was for original T). The attention scores from these different scales are then aggregated to form a "global" contextual representation. Let $(e_{ij}^{s})$ be the attention score between token $(x_{i})$ and $(x_{j})$ within its neighborhood $(\mathcal{N}_{i^{s}})$ at scale (s). The aggregated attention score considers information across scales:
$\alpha{ij}=\frac{\exp\left(\sum_{s}w_{s}\cdot e_{ij}^{s}\right)}{\sum_{k\in\bigcup_{s}\mathcal{N}i^{s}}\exp\left(\sum_{s}w_{s}\cdot e{ik}^{s}\right)}$ (1)
where$(w_{s})$ could be learnable weights or fixed, and the neighborhood $(\mathcal{N}_{i})$ spans relevant tokens across scales. We believe these design elements, ensure that the HSWA captures sufficient local details while effectively balancing the learning of both local and global information. The ablation studies presented in Section 4.3.1, especially regarding the performance of ACHDSWO (Acausal Hierarchical Dual Sliding Window with Overlap), further validate the effectiveness of this combined approach.
@@ 3.3 Boundary-aware Segmentation Loss
In temporal action segmentation, accurately capturing action transitions and segment-level dynamics across varying time scales is crucial. Surgical videos are typically long, and the duration of the segment fluctuates significantly, as shown in Table 1. To address this, we propose a Boundary-Aware Segmentation Loss ($\mathcal{L}_{\text{temporal}}$ ), a unified objective function that integrates multiple specialised loss components. To ensure precise action classification at each timestep and handle class imbalances prevalent in surgical datasets, we employ Focal Loss $\mathcal{L}_{\text{class}}$. To enhance the overlap between predicted and ground truth segments, we utilise Dice Loss derived from the Dice Coefficient:
@@
This formulation emphasises the overlap between predictions and ground truth, making it sensitive to both false positives and false negatives. While Focal and Dice losses address classification and overlap, they do not explicitly model the temporal smoothness of feature representations within an action segment. As illustrated in Figure 5, raw frame-to-frame feature similarities can be ambiguous. To encourage temporal consistency and smooth transitions, $\mathcal{L}_{\text{sim}}$ penalises deviations in feature representations $\mathbf{f}_{t}$ between consecutive frames. This penalty is weighted by a Gaussian function $G(t,t+1)$ which is designed to peak at the temporal centre of each ground truth action segment, diminishing towards its start and end points. This ensures that the strongest enforcement of feature similarity occurs within the core of an action, while allowing for more natural feature evolution near the transitions to adjacent actions, promoting coherent feature sequences.
$\mathcal{L}_{\text{sim}}=\sum_{t=1}^{T-1}G(t,t+1)\left(1-\cos(\mathbf{f}_{t},\mathbf{f}_{t+1})\right);G(t)=\exp\left(-\frac{(t-c)^{2}}{2\sigma^{2}}\right)$ (4)
Where $c$ is the temporal centre of the segment, $\mathbf{f}_{t}$ is the feature vector of frame $t$, and $\sigma$ is a hyperparameter controlling the width of the Gaussian weighting. This formulation ensures that the penalty for feature dissimilarity between consecutive frames ($\mathbf{f}_{t},\mathbf{f}_{t+1}$) is highest when those frames (referenced by index $t$) are near the centre ($c$) of their action segment and lowest when they are near the boundaries. This aligns perfectly with the goal of encouraging strong feature similarity at the "core of an action" while allowing for more natural feature evolution near the transitions.
To further refine boundary precision, we incorporate a Gaussian Similarity Truncated Mean Squared Error (MSE) loss. This loss focuses on boundary regions, applying a Gaussian weight to the MSE, which emphasises errors near segment boundaries:
@@
Where $G_{\text{boundary}}(t)$ is a Gaussian weight which is designed to have its highest values precisely at the annotated start and end frames of action segments (i.e., at the actual boundary points between differing actions), effectively creating high-importance zones for the loss calculation at these critical transitions. $\hat{b}_{t}$ and $b_{t}$ are the predicted and ground truth boundary positions at frame $t$, and $\tau$ is a truncation threshold to limit the influence of large errors. By integrating these loss components, the total temporal loss function is defined as:
@@
Where $\alpha$, $\beta$, $\gamma$, and $\delta$ = (1.0,0.2,0.5,0.5) were empirically determined on our validation set. This weighting factors balance the contributions of each loss term. This comprehensive loss function enables the model to effectively capture class imbalances, segment overlaps, temporal dependencies, and boundary precision, leading to improved performance in temporal action segmentation tasks.
@@ 4 Experiments and Analysis
Table 2: Comparison with state-of-the-art models Across Datasets (SAR-RARP, Cholec80, JIGSAW)
Dataset Model Acc Edit F1@10 F1@25 F1@50 SAR-RARP MS-TCNLi et al. (2020) 69.0 - 70.7 ActionCLIPPsychogyios et al. (2023) 80.4 80.6 MA-TCNVan Amsterdam et al. (2022) 83.4 81.6 81.7 ASRFIshikawa et al. (2021) $87.5\pm 3.7$ $83.6\pm 1.9$ $80.1\pm 3.2$ $70.8\pm 6.2$ $59.6\pm 3.2$ NETEYi et al. (2022) $87.8\pm 3.6$ $83.9\pm 1.5$ $78.7\pm 3.6$ $68.5\pm 3.2$ $54.6\pm 2.4$ AsformerYi et al. (2021) $\boldsymbol{88.8\pm 5.4}$ $\boldsymbol{84.1\pm 3.1}$ $\boldsymbol{84.3\pm 5.3}$ $68.1\pm 3.4$ $59.4\pm 3.1$ MSBATN (Ours) $88.7\pm 3.4$ $84.3\pm 2.7$ $84.0\pm 3.7$ $\boldsymbol{72.4\pm 2.6}$ $\boldsymbol{63.8\pm 1.5}$ Cholec80 ResNetYi et al. (2022) 78.3 57.8 52.2 PhaseLSTMNeil et al. (2016) 80.7 61.2 62.4 EndoNetTwinanda et al. (2016) 81.9 TeCNOCzempiel et al. (2020) 88.6 ASRFIshikawa et al. (2021) $82.5\pm 4.5$ $78.5\pm 3.5$ $84.3\pm 3.4$ $70.6\pm 5.7$ $60.1\pm 3.4$ AsformerYi et al. (2021) $82.8\pm 4.7$ $74.5\pm 6.7$ $84.1\pm 2.4$ $70.8\pm 6.2$ $57.1\pm 3.2$ NETEYi et al. (2022) $\boldsymbol{92.8\pm 5.0}$ $\boldsymbol{78.7\pm 9.4}$ $82.4\pm 3.4$ $70.1\pm 3.2$ $53.1\pm 4.8$ MSBATN (Ours) $89.4\pm 2.1$ $77.5\pm 8.0$ $\boldsymbol{84.4\pm 4.4}$ $\boldsymbol{73.1\pm 5.2}$ $\boldsymbol{60.2\pm 8.4}$ JIGSAW K-TCNVan Amsterdam et al. (2022) 83.8 86.3 90.4 MA-TCNVan Amsterdam et al. (2022) 86.8 90.1 93.6 ASRFIshikawa et al. (2021) $87.8\pm 4.5$ $89.0\pm 2.5$ $80.4\pm 6.5$ $70.1\pm 3.6$ $69.8\pm 2.5$ NETEYi et al. (2022) $90.8\pm 2.1$ $86.5\pm 3.5$ $74.4\pm 7.8$ $65.1\pm 3.7$ $58.6\pm 2.9$ AsformerYi et al. (2021) $\boldsymbol{92.8\pm 3.3}$ $\boldsymbol{89.8\pm 6.7}$ $83.5\pm 4.2$ $70.1\pm 3.5$ $65.4\pm 1.8$ MSBATN (Ours) $86.1\pm 6.3$ $89.5\pm 6.5$ $\boldsymbol{84.4\pm 5.5}$ $\boldsymbol{76.1\pm 3.1}$ $\boldsymbol{70.7\pm 2.8}$
Figure 6: Qualitative action segmentation comparison. (a) Ground Truth (GT), where each colour label represents an action. The rectangular boxes explicitly highlight actions for detailed comparison of actions’ start/end. (b) ASRF Ishikawa et al. (2021) and (c) Asformer Yi et al. (2021) show deviations in action start/end points (highlighted in red rectangular box). (d) The proposed MSBATN model demonstrates better alignment with GT (highlighted in green). Segments are zoomed from the full video timeline below.
The encoder ingests inputs of shape (T, 2048) and processes them through a dilated acausal TCN, effectively reducing the dimensionality to (T,256) while capturing temporal dependencies over varying scales. These embeddings are further refined through ten layers of multi-head sparse attention mechanisms (with 8 heads), incorporating advanced techniques such as sliding window attention for local dependencies and hierarchical attention for multi-scale temporal modelling. This combination enables the encoder to learn both fine-grained and long-range temporal patterns. The encoder produces frame-level predictions for action classes and boundary classes. To refine these initial predictions, three sequential decoders Zhang et al. (2023); Yi et al. (2021); Ishikawa et al. (2021); Yi et al. (2022) are employed. Each decoder shares a similar structure to the encoder but focuses on progressively refining predictions with higher accuracy and temporal smoothness. For optimisation, we utilised the Adam optimiser along with the softmax activation function to ensure stable training. In contrast to some existing approaches Yi et al. (2021), positional encoding was omitted, as the hierarchical and sliding attention mechanisms sufficiently capture temporal ordering and dependencies. Our boundary-aware loss function is designed to enhance the model’s segmentation and transition detection capabilities. It combines a weighted Dice loss, which penalises the model based on Intersection over Union (IoU), with a Gaussian Cosine Similarity loss that emphasises smooth transitions by modelling centre-to-boundary similarity. This composite loss function ensures robust learning of temporal dynamics and accurate identification of action boundaries. We split the dataset into 80-20 ratios for training testing for Jigsaw and Cholec80. We used dropout and early stopping to handle over-fitting and set the max epoch to 120. We have considered three different datasets such as SAR-RARP50 Psychogyios et al. (2023), JIGSAWS Gao et al. (2014), Cholec80 Twinanda et al. (2016), and for more details, please refer to supplementary material.
4.1 Evaluation metrics
$\text{Recall}=\frac{\text{TP}_{IoU>T}}{\text{TP}_{IoU>T}+\text{FN}_{IoU>T}}$
Compute Segmental F1 Score:
[…]
The Edit Score assesses the sequence of predicted action segments by comparing it to the ground truth sequence. It is calculated based on the Levenshtein distance, which quantifies the minimum number of edit operations (insertions, deletions, substitutions) required to transform the predicted sequence into the ground truth sequence. A higher Edit Score indicates a closer match between the predicted and actual sequences, reflecting the model’s proficiency in capturing the correct order and transitions of actions. This metric is less sensitive to minor temporal shifts but penalises over-segmentation and out-of-order predictions. Calculate the Levenshtein distance $D$ between the predicted and ground truth sequences. The Edit Score is then computed as Edit $\text{Edit Score}=1-\frac{D}{\max(L_{p},L_{g})}$ where $L_{p}$ and $L_{g}$ are the lengths of the predicted and ground truth sequences, respectively. The Segmental F1 Score is sensitive to over-segmentation and under-segmentation errors, penalising models that divide actions into too many segments or fewer segments. In contrast, the Edit Score penalises incorrect action order and over-segmentation but is more tolerant of slight temporal misalignments. The Segmental F1 Score evaluates the spatial and temporal precision of action boundaries, whereas the Edit Score assesses the structural and sequential accuracy of the predicted action sequence.
4.2 Results
Our Multi-staged Boundary-aware network has three main differences from Asformer. First, we employed acausal convolution layer in the initial layers of encoders and decoders. Secondly, we used hierarchical sparse multi-head window attention. Finally, we calculated centre sampled Boundary Aware Loss leverages the strengths of both cosine similarity and Gaussian functions, making it effective for segmentation tasks. As depicted in Table 2 our model outperforms other models in F1 scores and achieves comparable results in Acc and Edit with state of the art. While Accuracy and Edit Scores provide valuable insights into frame-level precision and strict sequential ordering, they are highly sensitive to slight temporal misalignment at segment boundaries. In contrast, the F1@k metric offers greater flexibility for boundary localisation (±$(\pm k\%)$ making it more robust to accurately segment long actions, even with minor shifts in timing. Figure 6 illustrates this qualitatively, showcasing MSBATN’s superior performance in accurately defining crucial action boundaries compared to other methods like ASRF and Asformer. Table 5 shows MSBATN is computationally efficient in comparison with other boundary aware model like ASRF Ishikawa et al. (2021).
Table 3: Comparison of different settings Across Datasets (SAR-RARP, Jigsaw, Cholec80)